# Image Can Bring Your Memory Back: A Novel Multi-Modal Guided Attack against Image Generation Model Unlearning

### v1

@@ 5.6.1. Generalizability
Table 4. Attack Success Rate (ASR, %) and Diversity Metrics (LPIPS, IS) with Different Reference Images.
Metric Method Task $R_{org}$ $R_{1}$ $R_{2}$ $R_{3}$
ASR(%) ESD Nudity 71.83 86.62 77.46 71.83
Church 96.00 94.00 96.00 92.00
UCE Nudity 76.76 77.46 75.35 78.24
Church 68.00 66.00 66.00 72.00
LPIPS ESD Nudity 0.42 0.40 0.44 0.41
Church 0.39 0.38 0.42 0.44
UCE Nudity 0.42 0.41 0.44 0.42
Church 0.37 0.38 0.42 0.44
IS ESD Nudity 4.36 4.20 4.50 4.42
Church 2.65 2.74 2.46 2.75
UCE Nudity 3.30 3.29 3.37 3.25
Church 2.72 2.75 2.75 2.94
These results confirm that RECALL is not simply copying or transferring a specific image, but is capable of robustly recalling erased content from a wide range of reference sources.
Generation Diversity.
To further investigate whether our method is fundamentally different from simple style-transfer or trivial image transformation, we quantitatively compare the diversity of generated images under three settings: image-only, text-only, and Recall (ours). The unlearning method used here is UCE, and the unlearning tasks include Nudity and Object-Church. As shown in Table 5, Recall achieves substantially higher diversity than image-only attacks and approaches the performance of text-only attacks (which do not use a reference image). Specifically, both LPIPS and IS for Recall consistently approach the upper bound, indicating that our attack does not simply copy or transform the reference image but instead encourages the model to recall the original content distribution associated with the target concept that should have been unlearned. Furthermore, the generated images presented in Appendix E.2, Figure 6, further illustrate that outputs from our Recall are highly diverse and distinctly different from the reference images.
Table 5. Diversity Comparison of Generated Images.
Metric Task Image-only Text-only RECALL
LPIPS Nudity 0.20 0.46 0.42
LPIPS Church 0.26 0.44 0.37
IS Nudity 3.30 4.77 4.36
IS Church 1.44 3.24 2.72
This confirms that RECALL leverages the model’s internal knowledge, rather than simply relying on the reference image, and provides evidence that the attack compromises true unlearning rather than performing surface-level style transfer.
Model Version Independence.
@@
Table 6. Attack Success Rate (ASR, %) on SD 2.x (UCE Unlearned) Across Four Tasks.
@@
5.6.2. Important Strategies and Parameters
Strategies
We analyze how different guidance strategies and key hyperparameters impact the effectiveness of the proposed Recall framework. Specifically, we evaluate various guidance modalities, including Text-only, Image-only, Text & R_noise, Text & Image, and our proposed Text & Adversarial Image. The detailed quantitative and visual results are presented in Sections 5.2 and 5.5, respectively. Empirical findings consistently indicate that combining textual prompts with adversarial image optimization significantly improves both attack performance and semantic consistency.
Furthermore, we demonstrate the benefits of the noise initialization strategy in terms of both the diversity and semantic fidelity of the generated images, as measured by LPIPS, IS, and CLIP Score. Table 7 reports the results on UCE-unlearned models across the Nudity and Church tasks. The results show that adopting noise initialization substantially improves diversity (higher LPIPS and IS) and semantic alignment (higher CLIP Score), confirming the effectiveness of this strategy in producing more meaningful and varied outputs. In addition, Appendix E (Table 9) confirms that periodically integrating $z_{ref}$ into $z_{adv}$ markedly boosts attack performance. Hence, we adopt this strategy throughout all experiments, setting $epoch_{\text{interval}}=5$ and $\gamma=0.05$ to maintain semantic consistency.
Table 7. Ablation results for noise initialization: comparison of diversity and semantic alignment metrics with (w/) and without (w/o) noise.
Task LPIPS IS CLIP Score
w/ w/o w/ w/o w/ w/o
Nudity 0.42 0.23 0.37 1.17 31.31 25.64
Church 4.36 3.11 2.72 1.52 27.82 23.75
Parameters
We further examine the sensitivity of attack performance to two critical optimization parameters:
Impact of Step Size ($\eta$). As shown in Appendix E Figure 5(a), reducing $\eta$ from 0.1 to 0.001 steadily improves the ASR, achieving optimal performance at $\eta=0.001$. Further reduction leads to diminished effectiveness due to insufficient gradient updates, indicating that $\eta=0.001$ offers the best balance.
Impact of Initial Balancing Factor ($\lambda$). Appendix E Figure 5(b) illustrates that increasing $\lambda$ from 0.10 to around 0.30 enhances the ASR before reaching saturation. Meanwhile, semantic alignment (CLIP score) peaks at $\lambda=0.25$, after which it declines, highlighting a trade-off between attack strength and semantic consistency. Consequently, $\lambda=0.25$ provides an optimal balance for effective attacks with high semantic fidelity.
6. Conclusion
In this paper, we propose Recall, a novel multi-modal adversarial framework explicitly designed to compromise the unlearning mechanisms of IGMs. Unlike prior text-based methods, Recall integrates adversarially optimized image prompts with fixed textual conditioning to induce unlearned IGMs to regenerate previously erased visual concepts. Extensive evaluations across multiple state-of-the-art unlearning techniques and four representative semantic scenarios reveal that current unlearning approaches remain vulnerable to our multi-modal attacks. These findings highlight the urgent need for more comprehensive and robust defense mechanisms to ensure the safety and trustworthiness of generative AI models.
Limitation and Future Work. Recall relies on semantically aligned reference images to guide the optimization, which may unintentionally introduce background biases. As a result, the optimized outputs can exhibit reduced diversity due to visual similarities with the reference image. While this has a limited impact on semantic fidelity or attack success, future work could focus on decoupling essential semantics from background artifacts during latent optimization to enhance image diversity and quality.
Appendix A Appendix Overview
This appendix presents supplementary material omitted from the main paper due to space constraints. Specifically, it includes:
  • Section B: Detailed experimental setup, covering datasets, unlearned IGMs, baseline methods, and evaluation metrics.
  • Section C: Complete algorithmic procedure of the proposed Recall framework.
  • Section D: Detailed configurations corresponding to the visual examples presented in the main paper.
  • Section E: Additional ablation studies on generalizability and hyperparameter sensitivity analysis.
Appendix B Experimental Setup
B.1. Datasets
We evaluate our method on four unlearning tasks: 1) Nudity, 2) Van Gogh-style, 3) Object-Church, and 4) Object-Parachute to ensure a thorough examination of unlearned models’ vulnerabilities. As multi-modality image generation needs the text and image at the same time, we first collect a reference image with the sensitive content by Flux-Uncensored-V2 (Flu, [n. d.]) (nudity, church, and parachute) and stable diffusion v2.1 (SDV, [n. d.]) (van Gogh) with a given text prompt for each task (as shown in Table 8), and then we adopted the text prompts used in UnlearnDiffAtk (Zhang et al., 2024) as the text prompts for each task, the details of these prompts are as follows:
  • Nudity: The Inappropriate Image Prompts (I2P) dataset (Schramowski et al., 2023) is involved, which contains a diverse set of prompts leading to unsafe or harmful content generation, including nudity. The number of prompts is 142.
  • Van Gogh-style: The involved prompts is artistic painting prompts introduce in ESD (Gandikota et al., 2023), the number of prompts is 50.
  • Object-Church and Object-Parachute: The prompts are generated by GPT-4 (OpenAI, 2023), and the number of prompts is 50 related to church and parachute, respectively.
Table 8. The reference image aligns with the text prompt and seed for the image-dependent attack method.
Task Nudity Van Gogh-style Object-Church Object-Parachute
Seed 6053 10710 33731 55254
B.2. Unlearned IGMs
The weights of involved unlearned SD models obtained by 10 state-of-the-art IGMU methods across four aforementioned unlearning tasks are sourced from three primary origins: ① the AdvUnlearn GitHub repository [^2] https://github.com/OPTML-Group/AdvUnlearn, as described in (Zhang et al., 2024); ② weights officially released by their respective authors, such as RECE (Gong et al., 2024), MACE (Lu et al., 2024) and DoCo (Wu et al., 2025); and ③ weights trained in-house using official implementations provided by ourselves.
B.3. Baselines
  • Text & Image: Both the original text prompt and a semantically relevant reference image containing the erased concept are provided as multi-modal inputs to the unlearned image generation models. This setting examines whether the reference image alone—without adversarial optimization—can facilitate the recovery of forgotten content and thereby expose the model’s residual memorization of the erased concept.
  • P4D (Chin et al., 2024): Prompting4Debugging (P4D) is a state-of-the-art attack that systematically discovers adversarial text prompts to bypass unlearned SD models. It leverages prompt optimization strategies to identify manipulations capable of eliciting forgotten concepts from the model. We report the results of P4D-K and P4D-N in this part simultaneously. We compare our method with P4D to demonstrate the advantages of adversarial image-based attacks over text-based adversarial prompting.
  • UnlearnDiffAtk (Zhang et al., 2024): UnlearnDiffAtk is a cutting-edge adversarial prompt generation technique tailored for evaluating unlearned diffusion models. It exploits the intrinsic classification properties of diffusion models with a given reference image to generate adversarial text prompts without requiring auxiliary classifiers or original SD models. We include this baseline to highlight the efficiency and effectiveness of our image-optimizing-based method in uncovering vulnerabilities in unlearned models.
P4D (with its variants P4D-K and P4D-N) (Chin et al., 2024) and UnlearnDiffAtk (Zhang et al., 2024) are text-optimization-based attack methods. And we use their officially released code (P4D [^3] https://github.com/joycenerd/P4D, UnlearnDiffAtk [^4] https://github.com/OPTML-Group/Diffusion-MU-Attack) with default configurations to extend the attack.
B.4. Evaluation Metrics
To evaluate the effectiveness of our proposed attack, we employ deep-learning-based detectors and classifiers tailored to each unlearning task. For the Nudity task, we adopt the NudeNet detector (Praneeth, 2023) with a detection threshold $\tau=0.45$ to identify explicit anatomical features. Following standard protocol (Ren et al., 2024), an image is considered a successful attack if it contains any of the following sensitive labels: MALE_BREAST_EXPOSED, MALE_GENITALIA_EXPOSED, FEMALE_BREAST_EXPOSED, FEMALE_GENITALIA_EXPOSED, BUTTOCKS_EXPOSED, or ANUS_EXPOSED. For the Van Gogh-style task, we use a Vision Transformer (ViT)-based (Dosovitskiy et al., 2021) style classifier pretrained on ImageNet and fine-tuned on the WikiArt dataset (Saleh and Elgammal, 2015), as in (Zhang et al., 2024), to verify whether the generated images exhibit Van Gogh’s characteristic artistic features. For the object-centric tasks—Object-Church and Object-Parachute—we leverage a ResNet-50 classifier pretrained on ImageNet to determine whether the corresponding object is present in the generated image.
Appendix C Algorithm
We list the Recall pipeline in Alg. 1, which could help readers to re-implement our method step-by-step.
Input: Reference image $P_{ref}$, randomly initialized image $P_{image}^{init}$, text prompt $P_{text}$, diffusion model $\mathcal{G}_{u}$ (with U-Net $\mathcal{F}_{\theta}$, text encoder $\mathcal{E}_{t}$, image encoder $\mathcal{E}_{i}$, image decoder $\mathcal{D}_{i}$), hyperparameters $\lambda$, $\gamma$, $\eta$, $\beta$, number of DDIM steps $T$, number of PGD iterations $N$.
1
Output: Adversarial image $P_{image}^{adv}$ for regenerating forgotten content.
2
[…]
4 $z_{ref}\leftarrow\mathcal{E}_{i}(P_{ref})$;
5 $z_{adv}\leftarrow\mathcal{E}_{i}(P_{image}^{init})$;
6 $h_{t}\leftarrow\mathcal{E}_{t}(P_{text})$;
7
[…]
9 for $t=T,T-1,\dots,1$ do
10       Compute noisy latents: $z_{\{ref,t}\},z_{\{adv,t\}}\leftarrow\sqrt{\bar{\alpha}_{t}}z_{\{ref,adv\}}+\sqrt{1-\bar{\alpha}_{t}}\epsilon_{t}$, $\epsilon_{t}\sim\mathcal{N}(0,I)$;
11
12      Predict noise residuals: $\hat{\epsilon}_{ref}\leftarrow\mathcal{F}_{\theta}(z_{\{ref,t\}},t,h_{t})$;
13       $\hat{\epsilon}_{adv}\leftarrow\mathcal{F}_{\theta}(z_{\{adv,t\}},t,h_{t})$;
14
15      Compute adversarial loss: $\mathcal{L}_{adv}\leftarrow\|\hat{\epsilon}_{ref}-\hat{\epsilon}_{adv}\|_{2}^{2}$;
16
17      $\nabla_{z_{adv}}\mathcal{L}_{adv}\leftarrow\text{Gradient of }\mathcal{L}_{adv}\text{ w.r.t }z_{adv}$;
18
19      for $i=1$ to $N$ do
20             $v_{i}=\beta\cdot v_{i-1}+\frac{\nabla_{z_{adv}}\mathcal{L}_{adv}}{\|\nabla_{z_{adv}}\mathcal{L}_{adv}\|_{1}+\omega}$;
21             $z_{adv}\leftarrow z_{adv}+\eta\cdot\text{sign}(v_{i})$;
22
23
24      if $t\mod epoch_{interval}==0$ then
25             $z_{adv}\leftarrow z_{adv}+\gamma\cdot z_{ref}$;
26
27
28
[…]
30
[…]
32
33return $I^{*}$;
Algorithm 1 Recall
Appendix D Details of Visualization Cases
To complement the qualitative results presented in Section 5.5, Table 10 provides detailed configurations used in generating the visual examples. This includes random seeds, guidance scales, and the corresponding input text prompts for various attack methods across four representative unlearning tasks: Nudity, Van Gogh-style, Object-Church, and Object-Parachute, evaluated on the unlearning models MACE (Lu et al., 2024) and RECE (Gong et al., 2024).
These details help interpret the outputs shown in Section 5.5, offering insight into how different attacks interact with unlearning constraints. Notably, baselines such as P4D (Chin et al., 2024) and UnlearnDiffAtk (Zhang et al., 2024) rely on heavily modifying the input text in order to bypass the unlearned models. While this occasionally restores erased content, it often degrades the semantic fidelity of the output image relative to the intended prompt—especially evident in the Nudity and Van Gogh-style cases.
In contrast, our Recall maintains the original prompt unchanged and leverages adversarial image guidance to effectively bypass unlearning while preserving strong semantic alignment. This distinction is clearly reflected in the table: where baselines yield incoherent or disjointed text inputs, Recall remains text-faithful yet adversarially potent.
Table 9. The attack performance of with (w/) and without (w/o) periodic integration.
Method Van Gogh-style Object-Church
w/ w/o w/ w/o
ESD 92.00 52.00 96.00 74.00
UCE 100.00 68.00 68.00 34.00
Table 10. Details of generating images under different attacks for MACE and RECE across different unlearning tasks.

### v2

@@ F.1 Reference Independence
Table 5: Attack Success Rate (ASR, %) and Diversity (LPIPS, IS, DINO) with Different Reference Images.
Method Ref. Nudity Object-Church
ASR $\uparrow$ LPIPS $\uparrow$ IS $\uparrow$ DINO $\uparrow$ ASR $\uparrow$ LPIPS $\uparrow$ IS $\uparrow$ DINO $\uparrow$
ESD $R_{\text{org}}$ 71.83 0.42 4.36 0.64 96.00 0.39 2.65 0.88
$R_{1}$ 86.62 0.40 4.20 0.61 94.00 0.38 2.74 0.89
$R_{2}$ 77.46 0.44 4.50 0.65 96.00 0.42 2.46 0.84
$R_{3}$ 71.83 0.41 4.42 0.62 92.00 0.44 2.75 0.89
UCE $R_{\text{org}}$ 76.76 0.42 3.30 0.69 68.00 0.37 2.72 0.92
$R_{1}$ 77.46 0.41 3.29 0.69 66.00 0.38 2.75 0.90
$R_{2}$ 75.35 0.44 3.37 0.70 66.00 0.42 2.75 0.92
$R_{3}$ 78.24 0.42 3.25 0.69 72.00 0.44 2.94 0.93
Figure 7: Randomly sampled images generated by the unlearned image generation model under our Recall attack, across four representative tasks. The visual results illustrate high diversity and semantic alignment with the text prompts, rather than mere reproduction of the reference images, confirming the effectiveness and generalizability of our approach.
F.2 Generation Diversity Across Methods
We assess whether Recall recovers a broader concept manifold, rather than reproducing a few memorized instances or performing trivial style transfer, by conducting a cross–method diversity comparison under two unlearning pipelines (ESD, UCE) and two tasks (Nudity, Object–Church). The evaluation includes weak baselines (Text-only, Image-only, Text&R_noise, Text&Image) and strong baselines (CCE, P4D, UnlearnDiffAtk, WACE) alongside Recall. Diversity is quantified with a DINO-based score. Results are shown in Figure 8.
Across both tasks and unlearning methods, Recall consistently exhibits higher diversity than Image-only, indicating that outputs do not collapse to copies or simple transforms of the reference image. It also surpasses Text&Image and Text&R_noise, suggesting that naïve multi-modal conditioning or noisy blending is insufficient to recover a broad concept manifold. Compared with the strong baselines, Recall reaches diversity that is competitive with, and often exceeds, CCE, P4D, UnlearnDiffAtk, and WACE; the trends are stable under both ESD and UCE, implying the advantage is not tied to a particular unlearning scheme.
Complementary qualitative evidence in Figure 7 shows randomly sampled outputs from Recall across four tasks. The results are visually diverse and non-homogeneous, rather than replications or near-duplicates of the reference images (see Table 3); they follow the semantics of the guiding text prompts while varying composition, layout, and appearance. Together with the DINO results in Figure 8, these observations indicate that Recall leverages the model’s internal concept space under joint text–image conditioning to recover a broader distribution of target-consistent samples, effectively addressing the concern on distributional coverage.
(a) Nudity
(b) Object-Church
Figure 8: Diversity across methods. DINO diversity scores for Nudity (left) and Object–Church (right) under ESD and UCE. Each category shows paired bars for ESD (blue) and UCE (orange). A vertical dashed line separates weak baselines (left) and strong baselines (right).
F.3 Model Version Independence
@@
Table 6: Attack Success Rate (ASR, %) on SD 2.x (UCE Unlearned) Across Four Tasks.
@@
Appendix G Ablation Study
Due to space constraints in the main paper, we present the ablation results for key strategies and hyperparameters in the adversarial optimization process here. Strategies include noise initialization and a periodic interval for injecting the reference latent $z_{ref}$ into the adversarial latent $z_{adv}$; hyperparameters include the step size ($\eta$) and the initial blending factor ($\lambda$).
G.1 Effect of Step Size $\eta$ on Attack Success Rate
Figure 9: Step size $\eta$ vs. attack success rate (ASR).
We first evaluate the influence of the step size $\eta$ on the attack success rate (ASR). As shown in Figure 9, ASR improves as $\eta$ decreases from 0.1 to 0.001, achieving peak performance around $\eta=0.001$. However, when $\eta$ is reduced further, the ASR begins to drop, likely due to insufficient gradient update magnitudes. This trend holds consistently across both ESD and UCE criteria, as well as across the Van Gogh and Church datasets, indicating that $\eta=0.001$ provides a balanced trade-off between stability and effectiveness.
G.2 Benefits of Initial Balancing on ASR, Semantic Alignment, and Efficiency
We study how the initial balancing factor $\lambda$, i.e., the proportion of reference features injected at initialization, affects the attack success rate (ASR), semantic alignment, and sampling steps. We sweep $\lambda\in[0.00,0.10,0.15,0.20,0.25,0.30,0.35,0.40,0.45,0.50]$ and report results in Figure 10.
In Figure 10(a), ASR increases with $\lambda$ and plateaus for $\lambda\gtrsim 0.30$. By contrast, the CLIP-based semantic alignment peaks near $\lambda\approx 0.25$ and then degrades as $\lambda$ grows, indicating that overly large injections bias the trajectory toward the reference branch and weaken text-conditioned alignment. Figure 10(b) further shows that moderate initialization ($\lambda\approx 0.20\!-\!0.30$) reduces the time consumption of attack required to succeed, yielding faster attacks without sacrificing semantic fidelity.
Overall, these trends delineate a practical trade space: larger $\lambda$ strengthens attack ability but can erode semantic consistency, whereas too small $\lambda$ slows convergence. An initial balancing factor around $\lambda=0.25$ provides a favorable operating point, shows high ASR, strong text alignment, and lower time cost.
(a) Balancing factor ($\lambda$) vs. ASR and CLIP Score.
(b) Balancing factor ($\lambda$) vs. ASR and Attack Time.
Figure 10: Ablation study of key hyperparameters: (a) effect of initial balancing factor $\lambda$ on ASR and semantic alignment (CLIP Score); (b) effect of initial balancing factor $\lambda$ on ASR and the needs of sampling step during the attacking.
G.3 Impact of Sampling Scheduler and Number of Steps
We examine how the sampling scheduler and the number of diffusion steps affect Recall’s performance. Specifically, we evaluate three schedulers, DDIM, LMS, and PNDM, under sampling steps $\{10,20,30,40,50\}$. Results are summarized in Figure 11.
As shown in Figure 11(a), the choice of scheduler has only a marginal effect on ASR (typically $1\%$ absolute difference across matched step counts). By contrast, the number of sampling steps has a more pronounced impact: ASR increases monotonically with additional steps but exhibits diminishing returns beyond 40 steps. We therefore adopt $50$ steps in the main experiments, which both attains the best observed ASR and aligns with the default configuration of Stable Diffusion v1.4, and facilitating fair comparison with prior work.
Figure 11(b) further reports the empirical density of the step index at which an attack first succeeds for two representative tasks (Nudity and Church) under two corresponding unlearning methods (ESD and UCE). The distributions concentrate on early steps, indicating that most successful attacks are completed well before the final step budget. These observations jointly suggest that Recall is largely scheduler-agnostic and achieves high ASR with moderate computational cost.
(a) ASR vs. sampling scheduler and number of steps.
(b) Empirical density of the step vs. succeed attack.
Figure 11: Ablation of sampling scheduler and number of steps. Panel (a) compares DDIM, LMS, and PNDM across step counts $\{10,20,30,40,50\}$. Panel (b) shows when attacks first succeed, indicating early-step concentration and thus computational efficiency.
G.4 Effect of Periodic Integration
To assess the benefit of periodically integrating the reference latent $z_{\text{ref}}$ into the adversarial latent $z_{\text{adv}}$, we study two key hyperparameters of periodic integration under two unlearned models (ESD, UCE) and two representative tasks (Nudity, Church). Specifically, we sweep the integration interval $epoch_{\text{interval}}\!\in\!\{1,2,4,5,10,20\}$ and the regularization coefficient $\gamma\!\in\![0,,0.10,0.15,0.20,0.25,0.30,0.35,0.40]$. Results are shown in Figure 12.
From Figure 12(a), moderate intervals (2255) consistently yield high ASR while keeping the required sampling steps low; very small intervals ($1$) can over-bias the trajectory toward the reference branch in some settings, whereas very large intervals ($\geq 10$) generally reduce ASR and increase step cost. From Figure 12(b), small-to-moderate regularization ($\gamma\approx 0.05$$0.20$) attains the best ASR; larger $\gamma$ gradually trades ASR for slightly higher DINO-based diversity, revealing a natural diversity–attack ability trade-off. Balancing effectiveness and efficiency across tasks and models, we adopt $epoch_{\text{interval}}=5$ and $\gamma=0.05$ in the main experiments.
(a) ASR (left axis, solid) and sampling step (right axis, dashed) vs. periodic interval ($\gamma=0.05$).
(b) ASR (left axis, solid) and diversity (right axis, dashed) vs. regularization coefficient ($\text{interval}=5$).
Figure 12: Ablation of periodic integration.
G.5 Sensitivity to Reference Alignment
We quantify how reference–image alignment influences attack performance. We add references from ImageNet and the open web that vary in semantic alignment to the nudity target: (i) matched ($R_{\text{org}}$); (ii) partially aligned (Bikini_man, Bikini_woman); and (iii) misaligned (Clothed, Bird); see Figure 13. With text prompts fixed, we evaluate two unlearning methods (ESD, UCE) and report attack success rate (ASR) and sampling steps to success (lower is better; larger values indicate success occurs later in the diffusion trajectory and hence lower efficiency); see Figure 14.
Across both ESD and UCE, alignment matters: partially aligned references yield higher ASR and fewer sampling steps than misaligned ones (Figure 14a–b). Within the partially aligned group, greater body exposure (Bikini_man) tends to improve ASR and reduce steps relative to Bikini_woman. When the reference is compositionally unrelated (Clothed, Bird), ASR drops and steps increase, approaching text-only behavior where occasional successes arise late in the trajectory and are largely attributable to the text prompt and stochastic sampling rather than the reference.
Recall remains functional with partially aligned references and degrades gracefully as alignment weakens, both in success rate and efficiency. When the reference is unrelated, performance approaches the text-only regime (lower ASR, higher steps), delineating the operational limits of reference guidance.
Figure 13: Reference-image cases for the nudity task. From left to right: matched reference $R_{\text{org}}$; partially aligned references Bikini_man and Bikini_woman; misaligned references Clothed and Bird. These cases vary primarily in semantic alignment to the target concept and are used to probe robustness and failure modes of reference guidance.
(a) ESD
(b) UCE
Figure 14: Sensitivity to reference alignment. Bars report ASR (left axis) and mean sampling steps to success (right axis) for five reference conditions: $R_{\text{org}}$ (matched), Bikini_man, Bikini_woman (partially aligned), Clothed, and Bird (misaligned). Partially aligned references improve success rate and efficiency relative to misaligned ones; unrelated references approach text-only behavior with lower ASR and higher required steps.
Table 7: Details of generating images under different attacks for MACE and RECE across different unlearning tasks.